OpenStudy (anonymous):

find a.b, when |a|=2,|b|=6 and the angle between a and b pi/4 radians

OpenStudy (anonymous):

a(vector).b(vector)=\[\left| a \right|\left| b \right| \cos \theta \]

OpenStudy (anonymous):

=\[2*6*\cos(\Pi/4)\] \[=12*(1/\sqrt{2})\]